x-1=y+1
x+1=2(y-1)

Answer by palanisamy(496)

The given equatins are,
x-1=y+1 ...(1)
x+1=2(y-1)...(2)
(1)-(2)=> -1-1 = y+1-[2y-2]
-2= y+1-2y+2
-2 = -y+3
y = 3+2
y = 5
Substituting in (1), x-1 = 5+1
x-1 = 6
x = 6+1
x = 7
Therefore the solution is (7,5)